# Foundation Models on a Budget: Approximating Blocks in Large Vision Models

### v5

@@ Abstract
Foundation Models have shown impressive performance in various tasks and domains, yet they require massive computational resources, raising concerns about accessibility and sustainability. Previous attempts to reduce foundation model size fall short of fully addressing the problem, as they end up increasing computational load through additional training steps. Recent works reveal that deep neural networks exhibit internal representation similarities. While inter-network similarities have enabled techniques such as model stitching and merging, intra-network similarities remain underexplored for improving efficiency. In this paper, we propose Transformer Blocks Approximation (TBA), a novel method that leverages intra-network similarities to identify and approximate transformer blocks in large vision models. TBA replaces these blocks using lightweight, closed-form transformations, without retraining or fine-tuning the rest of the model. The proposed method reduces the number of parameters while having minimal impact on the downstream task. We validate the effectiveness and generalizability of TBA through extensive experiments across multiple datasets (e.g., Imagenet-1k and CIFAR100) and state-of-the-art pretrained vision models (e.g, ViT, DiNO-v2, and DEiT).
@@ 1 Introduction
As Neural Networks (NNs) grow in size and complexity, their demand for computational resources has become a significant bottleneck. Despite the impressive performance of large models, they often come with substantial trade-offs, such as increased memory and power consumption. This has led to a growing interest in methods that can reduce model complexity without sacrificing performance. However, most existing approaches trying to mitigate these challenges either demand additional training steps or result in substantial performance degradation. Recent research showed that there exist similarities between representation within and between NNs. The observation that many layers or components within these networks perform similar functions or produce highly correlated outputs suggests that some blocks may be approximated, highlighting opportunities for network simplification. In this paper, we investigate two key research questions: (i) how to effectively identify blocks within neural networks that yield similar representations, and (ii) how to approximate these blocks with minimal impact on overall performance without any additional training steps. To address the first question, we provide an extensive analysis showing that the Mean Squared Error (MSE) effectively identifies such blocks, suggesting that their approximation should have a negligible impact. For the second question, we propose Transformer Blocks Approximation (TBA), a novel method that leverages similarities between representations to approximate transformer blocks using lightweight transformations, such as linear mappings. Once the blocks that have minimal impact on model functionality are identified, TBA replaces them with a simpler transformation. This approximation strategy allows for reducing the overall parameter count, ensuring minimal impact on downstream task performance. Our main contributions are as follows:
We introduce TBA, a general framework for identifying and approximating transformer blocks that produce similar outputs, using simpler transformations (e.g., linear), reducing model parameters and complexity with minimal to no impact on the downstream task performance (Figure 1).
@@ 2 Related work
Measuring similarities A range of metrics have been introduced to assess the similarity between latent spaces generated by different NNs (Klabunde et al., 2023; Ballester et al., 2023). One established approach is Canonical Correlation Analysis (CCA) (Hotelling, 1992), known for its invariance to linear transformations. Variants of CCA, such as Singular Value CCA (SVCCA) (Raghu et al., 2017), aim to enhance robustness, while techniques like Projection Weighted CCA (PWCCA) (Morcos et al., 2018) mitigate sensitivity to small perturbations. Another widely used metric, Centered Kernel Alignment (CKA) (Kornblith et al., 2019), captures the similarity between latent spaces while ignoring orthogonal transformations. However, recent work (Davari et al., 2022) highlights that this metric can be sensitive to shifts in the latent space. Additionally, Barannikov et al. (2021) proposes a method to compare two data representations by measuring the multi-scale topological dissimilarity, while Fumero et al. (2024) leverages the principles of spectral geometry to model and analyze the relationships between distinct latent spaces.
@@
Leveraging similarities Valeriani et al. (2024) examine the intrinsic dimensionality and neighbor compositions of representations in transformer models. Kvinge et al. (2022) investigate how models process variations in data points across layers, while Nguyen et al. (2020) assess the impact of network depth and width on hidden representations. Additionally, Crisostomi et al. (2023) study the conditions under which two latent spaces can be merged into a unified one. Moschella et al. (2023) construct a unified space shared by different NNs, enabling zero-shot stitching of independently trained models across different modalities (Norelli et al., 2023). More recently, Cannistraci et al. (2024) enable model stitching without explicit assumptions about the transformation class connecting the latent manifold embeddings, or with only partial correspondence between latent spaces (Cannistraci et al., 2023). Finally, Lähner & Moeller (2024); Maiorca et al. (2024) demonstrate that representations learned by distinct NNs can be aligned using simple transformations.
@@
Architectural efficiency While large-scale models with billions or even trillions of parameters continue to achieve state-of-the-art performance, their growth comes with trade-offs, such as slower inference times and significantly higher computational costs. Improving the efficiency of Deep Neural Network (DNN) has been a long-standing area of research. For instance, Veit et al. (2016) shows that removing residual blocks from deep Convolutional Neural Networks (CNNs) only marginally impacts performance, which inspired approaches to reduce inference time by dynamically deciding which layers to execute based on the input (Wu et al., 2018; Veit & Belongie, 2018). Additionally, various techniques to enhance efficiency have emerged, such as early exiting and model pruning. Early exit strategies, which introduce intermediate output layers at different stages of the network, have been shown to reduce inference time (Xin et al., 2020; Zhou et al., 2020; Yu et al., 2022; Tang et al., 2023). However, these approaches require the training of intermediate classifiers to enable exits at predefined layers. Alternatively, model pruning reduces computational load by either removing individual weights based on specific criteria, such as gradient information (Ma et al., 2023), entropy (Liao et al., 2023), or second-order information (Singh & Alistarh, 2020), or by eliminating larger structural components, like channels or residual blocks in ResNets (Bai et al., 2023; Wang & Wu, 2023) and self-attention layers in Transformers (Zhang & He, 2020; Sajjad et al., 2023; Venkataramanan et al., 2024; Zhang et al., 2024). Although effective, these approaches require training the model from scratch and, in the best case, fine-tuning. However, Bai et al. (2023) shows that for CNNs, this additional training step can sometimes be avoided.
@@
Unlike other methods, TBA leverages intra-network similarities to reduce foundation model complexity without the need for additional training steps while maintaining competitive performance.
3 Transformer Blocks Approximation
@@
The core idea of the proposed approach is to detect similar representations within NNs and replace the corresponding blocks with simple transformations. In the context of this work, a "block" refers to a sequence of layers including multi-head self-attention, normalization, and feed-forward layers, functioning as a cohesive unit. Instead of executing the entire NN, using transformer block approximation (TBA), we approximate these blocks, reducing computational complexity while preserving the network’s functionality. To provide a clear understanding, Figure 1 presents an overview of our method.
Identifying similar representations We hypothesize that certain foundation models may contain blocks that produce similar representations. To quantify these similarities, we employ the MSE. A low MSE between the output of an earlier block and a later block indicates that their respective representations are highly similar. This suggests that the intervening sequence of blocks contributes minimally to transforming the representation, highlighting potential redundancy.
Let represent the total number of blocks in the model. Let denote the output representation (a vector of dimension ) of block for a single input, where . For any two blocks, a starting block and an ending block (where ), we compute their output representations and . For each , where is a subset of points sampled uniformly at random from , defines the input to the corresponding layer . The MSE between these representations is defined as:
By systematically evaluating the MSE for various pairs of blocks , we can identify those being good candidates for approximation. This allows for targeted complexity reduction without substantially altering the network’s internal representations or downstream-task performance at a very low computational cost.
@@
Approximating transformer blocks Once two blocks and are identified as having highly similar output representations, our goal is to replace the intermediate blocks with a single, lightweight transformation that maps the output of block directly to an approximation of the output of block . This approach allows us to skip the computation of blocks , effectively reducing the overall computational costs. This approximation can be repeated for multiple, non-overlapping blocks, i.e., blocks and with .
@@
Let and represent the output representations from block and respectively, for the data points in . Our objective is to find a transformation such that:
@@
In this work, we consider to be a linear transformation that can be estimated by minimizing the squared error between the transformed output and the actual , which can be solved using least squares:
@@
This optimization problem allows for a closed-form solution that efficiently computes the optimal transformation . The solution bypasses the computation of all layers between any two blocks and , replacing them with . This approximation significantly reduces computational complexity by replacing one or more transformer blocks, comprising multi-head self-attention and feed-forward layers, with a low-cost linear transformation. To sum up, our approach consists of two main stages:
1.
Identification: We use MSE to identify blocks that produce similar representations.
2.
Approximation: We compute an efficient linear approximation using a transformation matrix to bypass the previously identified blocks.
This procedure reduces parameter count with minimal impact on internal representations, as shown in Figures 9, 7, 10, 11 and 8, and maintains compatibility with any downstream classifiers, achieving significant compression as shown Tables 1, 8, 9, 10, 11, 12 and 13.
@@ 4 Experiments
This section begins with an analysis of block-level similarities within pre-trained vision foundation models (Section 4.1), and we investigate how TBA impacts the representations of these latent spaces. Following this, Section 4.2 provides a comprehensive quantitative evaluation of TBA’s performance on various image classification tasks, including standard, zero-shot, and cross-dataset generalization scenarios. Lastly, in Section 4.3 we compare TBA against existing baselines, including an ablation study on transformation complexity, further validating the effectiveness of the proposed approach.
@@ 4.1 Latent analysis
Block similarities We investigate similarities emerging in latent representations using three distinct transformer-based models with different dimensionality: ViT-S, DiNO-B, and DEiT-S, and five datasets: CIFAR-10, CIFAR-100, MNIST, F-MNIST, and ImageNet1k.
Details on the models and datasets are provided in Table 6 and Table 7. Figure 2 presents the MSE matrices between blocks of different models when using different datasets. These similarities are calculated using only a small subset of the training data (i.e., 500 samples). Results reveal that while the patterns of similarity vary across models, they remain consistent across different datasets. This suggests that the similarity structure between computational blocks is predominantly influenced by the model rather than the dataset used. This finding aligns with observations from Nguyen et al. (2020), where DNNs trained from scratch tend to exhibit a distinctive “block structure” in their representations, linked to model overparameterization. Our results extend this observation to pretrained foundation models, where we show that such a structure is primarily influenced by the model. We conduct the analysis using the mean over the tokens, but in Section A.2.1 we show that this also applies when using only the CLS token from each block. Finally, in Figure 6, we present an ablation study to investigate the influence of the similarity metric. We correlate various metrics with downstream task performance using ViT-S and ImageNet1k. Our findings reveal that using simple metrics such as MSE or cosine similarity to identify block-level similarities exhibits a strong correlation with final accuracy. These results underscore the flexibility of our proposed method, indicating its compatibility with a range of similarity or distance functions that can capture inter-block relationships.
Effects of TBA on latent representations After identifying blocks producing similar representations, we analyze the impact on the latent representation output by the last block when approximating using TBA.
These blocks are approximated using a shared linear transformation applied across all tokens based on a subset of 3,000 training samples. For consistency, we employed the same models and datasets used for the results in Figure 2. We compute the MSE between the representations in the last block of the original and the TBA model when approximating the block. As illustrated in Figure 3, in most cases, the MSE increases as the block depth increases.
Original TBA

### v6

# : Transformer Optimization using Adaptive and Simple Transformations

@@ Abstract
Foundation models achieve State-of-the-art (SOTA) performance across different tasks, but their size and computational demands raise concerns about accessibility and sustainability. Existing efficiency methods often require additional retraining or fine-tuning, limiting their practicality. Recent findings suggest that deep neural networks exhibit internal representation similarities. While such similarities across different models have been exploited for enabling techniques such as model stitching and merging, intra-network redundancy remains underexplored as a source for efficiency gains. In this paper, we introduce Transformer Optimization using Adaptive and Simple Transformations (TOAST), a framework that exploits these redundancies to approximate entire transformer blocks with lightweight closed-form mappings, such as linear transformation or even the identity, without any additional training. Across SOTA pretrained vision models (e.g., ViT, DINOv2, DeiT) and datasets ranging from MNIST to ImageNet-1k, TOAST reduces parameters and computation while preserving, and in some cases improving, downstream performance. These results show that large portions of transformer depth can be replaced by trivial functions, opening a new perspective on efficient foundation models.
@@ 1 Introduction
As Neural Networks (NNs) continue to grow in size and complexity, their demand for computational resources has become a critical bottleneck. While larger models consistently achieve SOTA performance, this comes at the cost of substantial memory usage and power consumption, limiting their accessibility and deployment. This challenge is for instance most relevant in on-device scenarios, where saving memory, latency, and energy, even by little margins, is critical (Pan et al., 2022; Li et al., 2022). This has motivated a growing body of work on reducing model complexity. However, most existing approaches either require additional, resource-intensive training phases or lead to significant drops in accuracy. Recent studies reveal that there exists strong representational similarities both within and between NNs. In other words, when focusing on intra-network similarities, different blocks often perform overlapping functions or produce highly correlated outputs.
This redundancy suggests an opportunity: instead of retraining or pruning, can we approximate these blocks with simpler transformations? To address this question, we propose Transformer Optimization using Adaptive and Simple Transformations (TOAST), a novel framework that exploits block-level representational redundancy to replace transformer blocks with lightweight transformations. By doing so, TOAST reduces parameter count and computational cost, while maintaining (and in some cases even improving) downstream task performance. Crucially, our method is training-free, making it simple, efficient, and widely applicable, even in resource-constrained scenarios such as deployment on edge devices, where even the smallest available models may exceed memory or power budgets. Our main contributions are as follows:
We propose TOAST, a simple yet effective framework that replaces transformer blocks with lightweight transformations (e.g., linear maps or even the identity), significantly reducing parameters and computational cost while preserving downstream performance (Figure˜1).
@@ 2 Related work
A range of metrics have been introduced to assess the similarity between latent spaces generated by different NNs (Klabunde et al., 2023; Ballester et al., 2023). One established approach is Canonical Correlation Analysis (CCA) (Hotelling, 1992), known for its invariance to linear transformations. Variants of CCA, such as Singular Value CCA (SVCCA) (Raghu et al., 2017), aim to enhance robustness, while techniques like Projection Weighted CCA (PWCCA) (Morcos et al., 2018) mitigate sensitivity to small perturbations. Another widely used metric, Centered Kernel Alignment (CKA) (Kornblith et al., 2019), captures the similarity between latent spaces while ignoring orthogonal transformations. However, recent work (Davari et al., 2022) highlights that this metric can be sensitive to shifts in the latent space. Additionally, Barannikov et al. (2021) proposes a method to compare two data representations by measuring the multi-scale topological dissimilarity, while Fumero et al. (2024) leverages the principles of spectral geometry to model and analyze the relationships between distinct latent spaces.
Leveraging Similarities
@@
Valeriani et al. (2024) examine the intrinsic dimensionality and neighbor compositions of representations in transformer models. Kvinge et al. (2022) investigate how models process variations in data points across layers, while Nguyen et al. (2020) assess the impact of network depth and width on hidden representations. Additionally, Crisostomi et al. (2023) study the conditions under which two latent spaces can be merged into a unified one. Moschella et al. (2023) construct a unified space shared by different NNs, enabling zero-shot stitching of independently trained models across different modalities (Norelli et al., 2023). More recently, Cannistraci et al. (2024) enable model stitching without explicit assumptions about the transformation class connecting the latent manifold embeddings, or with only partial correspondence between latent spaces (Cannistraci et al., 2023). Finally, Lähner & Moeller (2024); Maiorca et al. (2024) demonstrate that representations learned by distinct NNs can be aligned using simple transformations.
Architectural Efficiency
@@
While large-scale models with billions or even trillions of parameters continue to achieve state-of-the-art performance, their growth comes with trade-offs, such as slower inference times and significantly higher computational costs. Improving the efficiency of Deep Neural Network (DNN) has been a long-standing area of research. For instance, Veit et al. (2016) shows that removing residual blocks from deep Convolutional Neural Networks (CNNs) only marginally impacts performance, which inspired approaches to reduce inference time by dynamically deciding which layers to execute based on the input (Wu et al., 2018; Veit & Belongie, 2018). Additionally, various techniques to enhance efficiency have emerged, such as early exiting and model pruning. Early exit strategies, which introduce intermediate output layers at different stages of the network, have been shown to reduce inference time (Xin et al., 2020; Zhou et al., 2020; Yu et al., 2022; Tang et al., 2023). However, these approaches require the training of intermediate classifiers to enable exits at predefined layers. Alternatively, model pruning reduces computational load by either removing individual weights based on specific criteria, such as gradient information (Ma et al., 2023), entropy (Liao et al., 2023), or second-order information (Singh & Alistarh, 2020), or by eliminating larger structural components, like channels or residual blocks in ResNets (Bai et al., 2023; Wang & Wu, 2023) and self-attention layers in Transformers (Zhang & He, 2020; Sajjad et al., 2023; Venkataramanan et al., 2024; Zhang et al., 2024). Although effective, these approaches require training the model from scratch and, in the best case, fine-tuning. However, Bai et al. (2023) shows that for CNNs, this additional training step can sometimes be avoided.
@@
Unlike other methods, TOAST leverages intra-network similarities to reduce vision transformers complexity without the need for additional training steps while maintaining competitive performance.
3 Blocks Approximation
@@
The central idea of our approach is that it is possible to leverage representation similarities within transformer-based architectures to replace entire attention blocks with simpler transformations. In this work, a block refers to a sequence of layers including multi-head self-attention, normalization, and feed-forward layers, that function together as a cohesive unit. By replacing these blocks with simpler transformations, we can reduce the computational complexity of the network while maintaining its core functionality.
Approximating Transformer Blocks
@@
Given two blocks and , our goal is to replace the intermediate blocks with a single, lightweight transformation that maps the output of block directly to an approximation of the output of block . This approach allows us to skip the computation of blocks , effectively reducing the overall computational costs. This approximation can be repeated for multiple, non-overlapping blocks, i.e., blocks and with . An overview of the method is provided in Figure˜1.
@@
Let and represent the output representations from block and respectively, for the data points in , sampled uniformly at random from the full training dataset . Our objective is to find a transformation such that:
@@
In this work, we consider to be the identity or a linear transformation . We can compute the linear transformation by minimizing the squared error between the transformed output and the actual :
@@
This optimization problem allows for a closed-form solution that efficiently computes the optimal transformation . The solution bypasses the computation of all layers between any two blocks and , replacing them with . This approximation reduces computational complexity while minimally affecting internal representations, as illustrated in Figures˜9, 7, 10, 11 and 8, and preserves compatibility with downstream classifiers, achieving significant compression as shown in Tables˜1, 2, 3, 7, 8, 9 and 10.
Patterns of Similarity between Transformer Blocks
Inspired by existing results Kornblith et al. (2019); Nguyen et al. (2020), and in particular Venkataramanan et al. (2024), which show that multi-head attention modules exhibit similarity in learned representations, we investigate whether pre-trained foundation models contain entire blocks that produce highly similar representations. To investigate this hypothesis, we adopt the CKA, where a high score between the outputs of two blocks indicates that their respective representations are highly similar. This, in turn, suggests that the intermediate blocks contribute minimally to transforming the representation, highlighting potential redundancy in the model.
After centering the representations, i.e., and , by subtracting their respective means, the CKA is computed as:
Here, denotes the Frobenius norm. A CKA value closer to 1 indicates a higher degree of similarity between the representations of the two blocks.
By systematically evaluating the CKA across pairs of blocks , it is possible to distinguish patterns of similarities between different blocks. These patterns indicate opportunities for targeted complexity reduction, enabling the approximation of redundant blocks without substantially altering internal representations or downstream task performance.
@@ 4 Experiments
In this section, we first analyze the similarities between different transformer blocks to motivate their approximation using simple transformations. We then present comprehensive results on image classification across various models and datasets to demonstrate the effectiveness and efficiency of the proposed method. Beyond these core results, we further study the robustness of TOAST through ablations on the number of samples required for approximation and the choice of translator architecture. Overall, our findings show that TOAST achieves strong performance while producing lighter and faster models. Due to space constraints, additional results on zero-shot image classification, as well as further qualitative and quantitative analyses, are provided in the Appendix (Sections˜A.2.3 and A.2.2).
@@ 4.1 Latent analysis
In this section we investigate similarities in the latent representations of DiNO-B and DEiT-S on five datasets: CIFAR-10, CIFAR-100, MNIST, F-MNIST, and ImageNet1k. We compute the CKA similarity metric averaged over all tokens from a small subset of 100 training samples, which is sufficient to reveal block-level similarity patterns while remaining computationally efficient. Additional results with other pretrained vision transformers (ViT-T, ViT-S, DiNO-S, ViT-B) are provided in Section˜A.2.1, showing consistent patterns for each model across different datasets.
Do vision transformer models exhibit block-wise similarity patterns?
The results in Figure˜2 reveal that while the similarity patterns differ across models, they remain largely consistent for the same model across different datasets. This suggests that the similarity structure between computational blocks is predominantly influenced by the model itself. Although the general similarity pattern remains the same, the differences in CKA values become more pronounced (i.e., the block structure becomes more evident) as the complexity of the dataset increases (e.g., from MNIST to ImageNet1k). These finding aligns with observations from Nguyen et al. (2020), where DNN trained from scratch exhibit a distinctive "block structure" in their representations, which is linked to model overparameterization. Our results extend this observation to vision pre-trained foundation models, showing that such a structure is primarily an intrinsic property of the model. Moreover, these consistent block-wise patterns indicate potential targets for approximation, suggesting that entire blocks may be replaced with simpler transformations without substantially altering the model’s internal representations.
How does TOAST affect latent representations?
We next analyze the impact of the proposed transformations on the final block’s latent representations, which are used for downstream classification. We approximate these blocks using a shared linear transformation applied across all tokens, estimated on a subset of 500 training samples. For consistency, we use the same models and datasets as in Figure˜2.
Original
TOAST